# Abundance of cosmic voids in EFT of dark energy

### v1

@@ V Conclusions
In addition to the above qualitative trends, we have noted that there are theoretical bounds on the parameters coming from the requirement that the scalar field must take a real value during the course of the void formation. For example, this requirement gives $p\gtrsim 0.85$ for $\alpha_{B0}=0.1$.
@@
In this paper, we have focused on the impacts of the EFT parameters characterizing Horndeski gravity. It would be interesting to extend the present analysis to more general theories of gravity, such as degenerate higher-order scalar-tensor theories [Langlois:2015cwa], by incorporating additional EFT parameters. It is fair to say that many aspects of actual complex process of the void formation are ignored in the simple spherical model, and hence our study is a first step toward understanding in more detail the formation of voids in the EFT of dark energy and modified gravity. To this end, it is important to perform $N$-body simulations. Nevertheless, we believe that the simplified treatment with spherical symmetry presented in this paper helps us to understand, at least qualitatively, how the void formation is modified in the EFT of dark energy.
@@ Appendix A Relation between the Horndeski theory and the EFT of dark energy
The action for the Horndeski theory [Horndeski:1974wa] up to quadratic terms of the second derivatives of the scalar field $\phi$ is given by [Deffayet:2011gz, Kobayashi:2011nu]
@@
The $\alpha$-basis formulation of the EFT of dark energy is particularly useful for comparing theoretical predictions with observational data. Following Refs. [Bellini:2014fua, Cusin:2017mzw] (see also [Gleyzes:2014rba, Langlois:2017mxy, Dima:2017pwp]), one can relate the above Horndeski action to the EFT.
@@
where $H:=\dot{a}/a$, $\delta N:=N-1$, $\delta K^{j}_{i}:=K^{j}_{i}-H\delta^{j}_{i}$, and $\delta\mathcal{K}_{2}:=\delta K^{2}-\delta K^{j}_{i}\delta K^{i}_{j}$. We have kept higher-order terms that are responsible for the Vainshtein mechanism effective on small scales. The relation between the coefficients in the above action and the functions appearing in the Horndeski action is given by [^1] Different conventions are adopted in the literature for the definition of $\alpha_{B}$ and the sign in front of $G_{3}$ in the action. Our convention coincides with that in Refs. [Bellini:2014fua, Zumalacarregui:2016pph], but our $\alpha_{B}$ differs by a factor of 2-2 from that in Ref. [Dima:2017pwp].
@@ Appendix C Basic equation for the spherical collapse model
To compute the void size function in Sec. IV, we also need to evaluate the critical density contrast $\delta_{\mathrm{c}}$ for halo formation. For this purpose, we employ the spherical collapse model, which is widely used to estimate $\delta_{\mathrm{c}}$ semi-analytically [Bellini:2012qn, Barreira:2013xea, Albuquerque:2024hwv, Takadera:2025oyf]. In the following, we present the basic equations governing the evolution of a top-hat matter overdensity.
@@ Appendix D Excursion set theory and Sheth–van de Weygaert void size function
In this appendix, we briefly review the excursion set theory [Bond:1990iw] and derive the Sheth–van de Weygaert void size function [Sheth:2003py] following Ref. [Zentner:2006vw].
@@
This result was derived in Ref. [Sheth:2003py]. In the main text, we use $\mathrm{d}F/\mathrm{d}\sigma=2\mathrm{d}F/\mathrm{d}S$ rather than $\mathrm{d}F/\mathrm{d}S$ as the differential probability, and thus [Jennings:2013nsa]

### v2

@@ V Conclusions
In addition to the above qualitative trends, we have noted that there are theoretical bounds on the parameters coming from the requirement that the scalar field must take a real value during the course of the void formation. For example, this requirement gives $p\gtrsim 0.87$ for $\alpha_{B0}=0.1$.
@@
In this paper, we have focused on the impacts of the EFT parameters characterizing Horndeski gravity. It would be interesting to extend the present analysis to more general theories of gravity, such as degenerate higher-order scalar-tensor theories [33], by incorporating additional EFT parameters. It is fair to say that many aspects of actual complex process of the void formation are ignored in the simple spherical model, and hence our study is a first step toward understanding in more detail the formation of voids in the EFT of dark energy and modified gravity. To this end, it is important to perform $N$-body simulations. Nevertheless, we believe that the simplified treatment with spherical symmetry presented in this paper helps us to understand, at least qualitatively, how the void formation is modified in the EFT of dark energy.
@@ Appendix A Relation between the Horndeski theory and the EFT of dark energy
The action for the Horndeski theory [25] up to quadratic terms of the second derivatives of the scalar field $\phi$ is given by [17, 29]
@@
The $\alpha$-basis formulation of the EFT of dark energy is particularly useful for comparing theoretical predictions with observational data. Following Refs. [7, 14] (see also [23, 32, 19]), one can relate the above Horndeski action to the EFT.
@@
where $H:=\dot{a}/a$, $\delta N:=N-1$, $\delta K^{j}_{i}:=K^{j}_{i}-H\delta^{j}_{i}$, and $\delta\mathcal{K}_{2}:=\delta K^{2}-\delta K^{j}_{i}\delta K^{i}_{j}$. We have kept higher-order terms that are responsible for the Vainshtein mechanism effective on small scales. The relation between the coefficients in the above action and the functions appearing in the Horndeski action is given by [^1] Different conventions are adopted in the literature for the definition of $\alpha_{B}$ and the sign in front of $G_{3}$ in the action. Our convention coincides with that in Refs. [7, 53], but our $\alpha_{B}$ differs by a factor of 2-2 from that in Ref. [19].
@@ Appendix C Basic equation for the spherical collapse model
To compute the void size function in Sec. IV, we also need to evaluate the critical density contrast $\delta_{\mathrm{c}}$ for halo formation. For this purpose, we employ the spherical collapse model, which is widely used to estimate $\delta_{\mathrm{c}}$ semi-analytically [6, 5, 1, 46]. In the following, we present the basic equations governing the evolution of a top-hat matter overdensity.
@@ Appendix D Excursion set theory and Sheth–van de Weygaert void size function
In this appendix, we briefly review the excursion set theory [9] and derive the Sheth–van de Weygaert void size function [44] following Ref. [51].
@@
This result was derived in Ref. [44]. In the main text, we use $\mathrm{d}F/\mathrm{d}\sigma=2\mathrm{d}F/\mathrm{d}S$ rather than $\mathrm{d}F/\mathrm{d}S$ as the differential probability, and thus [26]